OpenStudy (anonymous):

I think I have the right answer, but I am not sure. Solve the system by the elimination method. x + y - 6 = 0 x - y - 8 = 0 When you eliminate y , what is the resulting equation?

OpenStudy (anonymous):

2x=14

OpenStudy (babylove18):

yes your right

OpenStudy (phi):

they may want you to write it as 2x-14=0 though it is true 2x=14 and also x= 7

OpenStudy (anonymous):

great, thanks!

OpenStudy (phi):

notice you add the two equations and that is what you get 2x + 0 -14= 0 which we would write as 2x-14=0

OpenStudy (anonymous):

Can you help with another?

OpenStudy (phi):

did you post it?

OpenStudy (anonymous):

Solve the system by the elimination method. 3x - 2y - 7 = 0 5x + y - 3 = 0 To eliminate y, the LCM is 2. Which of the following is the resulting equations?

OpenStudy (anonymous):

1. 3x - 2y - 7 = 0 5x + y - 3 = 0 - 2. 3x - 2y - 7 = 0 -10x - 2y + 6 = 0 - 3. 3x - 2y - 7 = 0 10x + 2y - 6 = 0

OpenStudy (anonymous):

My question is, do you multiply each term by 2? Or just "y"?

OpenStudy (phi):

you multiply both sides of the equation by 2 (both sides to keep it equal) think of it as: double the left side and double the right side of course, the right side is 0 and 2*0 is still 0 on the left side, you have to double (i.e. multiply by 2) each term

OpenStudy (phi):

it's easier to see with numbers: 1+1 = 2 if you multiply both sides by 2 you write: 2*(1+1) = 2*2 2*1 + 2*1 = 4 2+2= 4 4=4 it works

OpenStudy (anonymous):

So 2* (3x) (- 2y) (- 7) = 0 (2) 6x-4y-14=0 And 5x + y - 3 = 0 10x+2y-6=0

OpenStudy (phi):

yes, you could multiply the top eq. by 2 but we don't want to do that. we want the number in front of the y to be the same number. ( or the minus of it). so we only do the bottom equation.

OpenStudy (anonymous):

So the answer would be 3. 3x - 2y - 7 = 0 10x + 2y - 6 = 0?

OpenStudy (phi):

yes. the idea is both equations are still true (multiplying the 2nd eq by 2 does not make it wrong, just different) now we add them we would get 13x +0 - 13=0 or just 13x -13=0 and we are are our way to solving for x

OpenStudy (anonymous):

Thank you.